PLEASE POST EXCEL SPREADSHEET

  1. Please show all computation and underline your final answer in your submission.

Below are several financial statement items for fiscal year 2016 for two grocery chains, Whole Foods Market, an upscale organic grocer, and The Kroger Co. a mainstream grocer. ($ millions)

($ millions)

Whole Foods Market

The Kroger Co.

Net income

$     507

$    1,957

Sales

15,724

115,337

Average assets

6,041

35,201

Average stockholders’ equity

3,497

6,754

Calculate each company’s return on assets (ROA) and return on equity (ROE). Comment on any differences you observe.

Disaggregate the ROA for each company into profit margin (PM) and asset turnover (AT). Explain why Whole Foods has a higher ROA, is it because of PM or AT or both?
Answer #1
$ millions) Whole Foods Market The Kroger Co.
i Net income                 507               1,957
ii Sales            15,724           115,337
iii Average assets              6,041             35,201
iv Average stockholders’ equity              3,497               6,754
answer a =i/iii ROA = Net income/Average asset 8.39% 5.56%
answer b =i/iv ROE = Net income/Average equity 14.50% 28.98%
ans -c we can see that whole food has higher ROA but lower ROE. Kroger co is able to make more share holder wealth compared to Whole food
Kroger is better company from share holder point of view
ans -d ROA disaggregation
ROA = Net income/Total asset = (Net income/Sales) *(Sales/Total asset)= Profit margin* Asset turnover
Profit margin =i/ii 3.2% 1.7%
Asset turnover = ii/iii                    2.60                     3.28
ROA for Whole food is higher due to higher profit margin . Asset turnover is lower compared to Kroger